OpenStudy (anonymous):

How do you do Pythagorean Theorem questions that involve x * the square root of y?

OpenStudy (alexalas):

Can you give me an example of this problem

OpenStudy (anonymous):

@alexalas An example would be for the triangle a = 3 and b = 8 (this is an example) and some how it has to fit the equation that I gave.

OpenStudy (anonymous):

My actual problem resides in the confines of a = 5 and b = 4

OpenStudy (alexalas):

Do you have a picture of the problem. I work better with pictures.

OpenStudy (anonymous):

Let me screen shot and send it right quick

OpenStudy (anonymous):

@alexalas here

OpenStudy (alexalas):

Okay here is what you do. a^2+b^2=c^2 what is a^2 + b^2

OpenStudy (anonymous):

c = 41

OpenStudy (anonymous):

The square root of 41 = 6.4...

OpenStudy (alexalas):

41 does not square root into a whole number

OpenStudy (alexalas):

take \[\sqrt{125}\] as an example you factor it into \[\sqrt{25*5}\] and then square root 25 and it will look like \[5\sqrt{5}\] but 41 is a prime number so you can't factor out any numbers. so the answer would just be(you aswer)

OpenStudy (alexalas):

you there?

OpenStudy (anonymous):

It is just REALLY confusing and hard to understand.